# I Predict Therefore I Am: Is Next Token Prediction Enough to Learn Human-Interpretable Concepts from Data?

### v1

@@ 4.3 Unifying the Linear Representation Hypothesis
Corollary 4.2 reveals that the weights $\mathbf{W}$ learned through linear probing are intimately connected to the linear transformation $\mathbf{A}$ from our identifiability results, with the key relationship $\mathbf{W}\mathbf{A}=\mathbf{I}$. This relationship, previously unexplored, provides a unified framework for understanding the Linear Representation Hypothesis across various manifestations.
These manifestations include Concepts as Directions, Concept Manipulability, and Linear Probing, all of which are unified through the linear transformation $\mathbf{A}$. Specifically:
  • Concepts as directions and concept manipulability focus on extracting or manipulating a columns of $\mathbf{A}$, which correspond to the latent concepts of interest.
@@
By leveraging the relationship $\mathbf{W}\mathbf{A}=\mathbf{I}$, we demonstrate that these different forms of the linear representation hypothesis are not independent, but are instead intrinsically linked through the same linear transformation $\mathbf{A}$ from our identifiability result. This unified viewpoint underscores how $\mathbf{A}$ serves as the underlying structure that connects concept representation, manipulation, and probing.
@@ 5 Experiments on Simulation and Real Data
We begin by conducting experiments on synthetic data, which is generated through the following process: First, we create random directed acyclic graphs (DAGs) with $n$ latent variables, representing concepts. For each random DAG, the conditional probabilities of each variable given its parents are modeled using Bernoulli distributions, where the parameters are sampled uniformly from $[0.2,0.8]$. To simulate a nonlinear mixture process, we then convert the latent variable samples into one-hot format and randomly apply a permutation matrix to the one-hot encoding, generating one-hot observed samples. These are then transformed into binary observed samples. To simulate next-token prediction, we randomly mask part of the binary observed data, i.e., ${x}_{i}$, and use the remaining portion, i.e., $\mathbf{x}_{j}$ where $j\neq i$, to predict the masked part ${x}_{i}$. More details can be found in Section G.1 in Appendix.
@@
In Theorem 3.1, we demonstrate that the features learned through next-token prediction approximate a linear transformation of $\log p(\mathbf{c}|\mathbf{x})$. This approximation becomes tighter when the mapping from $\mathbf{c}$ to $\mathbf{x}$ is approximately invertible, as analyzed in Section 3.2. Building on this, Corollary 4.2 establishes that for a data pair $(\mathbf{c},\mathbf{x})$, the representations $\mathbf{f}_{\mathbf{x}}(\mathbf{x})$ is linearly separable. Therefore, to evaluate the linear transformation described in Theorem 3.1, we assess the degree to which the learned features can be classified linearly for data pairs $(\mathbf{c},\mathbf{x})$.
Figure 2: (Left) Classification accuracy improves with the size of observed variables, consistent with the approximation tightness analysis in Section 3.2. (Right) Classification accuracy across different latent graph structures (ER1, ER2, ER3) demonstrates the robustness of the identifiability results to graph structure and latent variable size.
@@
Figure 3: Results of the product $\mathbf{A}_{s}\times\mathbf{W}_{s}$ across the LLaMA-2 and Pythia model families. Here, $\mathbf{A}_{s}$ represents a matrix derived from the feature differences of 27 counterfactual pairs, while $\mathbf{W}_{s}$ is a weight matrix obtained from a linear classifier trained on these features. The product approximates the identity matrix, supporting the theoretical findings outlined in Corollary 4.2. Refer to Appendix I for more results on LLaMA-3 and DeepSeek-R1.
@@
We finally present experiments on pre-trained LLMs. While it is challenging to collect all latent variables from real-world data to directly evaluate our identifiability results, we can instead assess the corollaries. For Corollary 4.1, prior studies have already demonstrated the linear representation properties of LLM embeddings using counterfactual pairs or linear probing (Mikolov et al., 2013; Pennington et al., 2014; Turner et al., 2023; Li et al., 2024; Wang et al., 2023; Park et al., 2023). Building on this, we shift our focus to a new property in Corollary 4.2: the weight matrix $\mathbf{W}$ is an invertible matrix related to the linear transformation matrix $\mathbf{A}$. This property has not been explored in previous works. To investigate it, we utilize 27 counterfactual pairs from Park et al., which are constructed based on the Big Analogy Test dataset (Gladkova et al., 2016). More details can be found in Section G.2 in Appendix. Specifically, supported by Corollary 4.1, we compute the differences between the features of these 27 counterfactual pairs on pre-trained LLMs to construct a matrix $\mathbf{A}_{s}$ with dimensions $27\times dim$, where $dim$ corresponds to the feature dimension of the pre-trained large language model (LLM) used. This matrix can be regarded as a submatrix of $\mathbf{A}$, capturing the pairwise feature differences between the 27 counterfactual pairs. Additionally, we train a linear classifier using the features of these counterfactual pairs and extract the corresponding weights, supported by Corollary 4.2, forming a weight matrix $\mathbf{W}_{s}$ with dimensions $dim\times 27$, which also serves as a submatrix of $\mathbf{W}$. Finally, we examine the product of $\mathbf{A}_{s}$ and $\mathbf{W}_{s}$ to assess whether it approximates the identity matrix. Figure 3 displays the results of this product across the LLaMA-2 and Pythia model families. The results show that the product approximates the identity matrix, which is consistent with the theoretical finding in Corollary 4.2.
@@ 6 Conclusion and Discussion
Our analysis of LLMs through the lens of latent variable models and identifiability provides valuable insights into the underlying mechanisms that contribute to their success [^2] We strongly encourage the reader to refer to the additional discussions on latent variable models in Appendix H.. We have shown that that the representations learned by LLMs can be approximated as a linear transformation of the posterior distribution of latent variables. This not only reinforces the linear representation hypothesis but also highlights that concepts are encoded in the linear transformation matrix.
@@
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Related Work
This work is closely related to causal representation learning (Schölkopf et al., 2021), which aims to identify high-level latent causal variables from low-level observational data. Many prior studies (Brehmer et al., 2022; Von Kügelgen et al., 2021; Massidda et al., 2023; von Kügelgen et al., 2023; Ahuja et al., 2023; Seigal et al., 2022; Shen et al., 2022; Liu et al., 2022a; Buchholz et al., 2023; Varici et al., 2023; Liu et al., 2024b, 2022b, a, c; Hyvarinen & Morioka, 2016; Hyvarinen et al., 2019; Khemakhem et al., 2020) have developed theoretical frameworks supporting the recovery of true latent variables up to simple transformations. However, these works primarily focus on continuous spaces and do not address the next-token prediction framework employed by LLMs. A subset of works (Gu & Dunson, 2023; Kong et al., 2024; Kivva et al., 2021) has explored causal representation learning in discrete spaces for both latent and observed variables, often imposing specific graph structures and assuming invertible mappings from latent to observed spaces. Despite their focus on discrete settings, none of these studies consider the next-token prediction framework in LLMs. In contrast, our work overcomes these limitations. We analyze approximate identifiability without relying on strict invertibility assumptions, which better aligns with the complex and often non-invertible relationships observed in real-world data. While a recent study has examined non-invertible mappings from latent to observed spaces, they rely on additional historical information to effectively restore invertibility. In our approach, we focus on achieving approximate identifiability without requiring such additional information. A very recent work (Rajendran et al., 2024) explores identifiability analysis for LLMs; however, they model both observed text data and latent variables as continuous and still assume invertibility in their analysis. In contrast, our framework explicitly considers discrete latent and observed variables, relaxed invertibility assumptions, and directly aligns with the next-token prediction paradigm, offering a more realistic and generalizable approach to causal representation learning in LLMs.
@@ Appendix B Limitations
We then use Jensen’s inequality to approximate the right of Eq. (10) as follows:
$\log\sum_{\mathbf{c}}{p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x})}=\log\sum_{\mathbf{c}}p(\mathbf{c}|{y})\frac{p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x})}{p(\mathbf{c}|{y})}\geq\sum_{\mathbf{c}}p(\mathbf{c}|{y})\log\frac{p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x})}{p(\mathbf{c}|{y})}.$ (11)
Expanding the right-hand side of Eq. (11), we obtain:
$\log\sum_{\mathbf{c}}p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x})\geq\sum_{\mathbf{c}}p(\mathbf{c}|y)\log p(\mathbf{c}|\mathbf{x})+\sum_{\mathbf{c}}p(\mathbf{c}|y)\log p(y|\mathbf{c})-\sum_{\mathbf{c}}p(\mathbf{c}|y)\log p(\mathbf{c}|y).$ (12)
We now analyze the three terms on the right-hand side of Eq. (12) in more detail. For the first term $\sum_{\mathbf{c}}p(\mathbf{c}|y)\log p(\mathbf{c}|\mathbf{x})$, this represents a weighted average of $\log p(\mathbf{c}|\mathbf{x})$, reflecting the contribution of $p(\mathbf{c}|\mathbf{x})$ over all $\mathbf{c}$. In the context in this work, $\mathbf{c}$ represents discrete variables. If $\mathbf{c}$ is assumed to have a finite set of enumerated values, the summation is implicitly understood within $p(\mathbf{c}|y)\log p(\mathbf{c}|\mathbf{x})$, eliminating the need to explicitly include $\sum_{\mathbf{c}}$. Consequently, the first term can be written more succinctly as $p(\mathbf{c}|y)\log p(\mathbf{c}|\mathbf{x})$. For the last two terms, both involve a summation over discrete variables $\mathbf{c}$ and can therefore be treated as independent of $\mathbf{c}$. As a result, they can be regarded as constants independent of $\mathbf{c}$. With these simplifications, we arrive at the following approximate form:
$\log\big{(}\sum_{c}p(y|\mathbf{c})p(\mathbf{c}|\mathbf{x})\big{)}\approx p(\mathbf{c}|y)\log p(\mathbf{c}|\mathbf{x})+\mathbf{b}_{y},$ (13)
where $\mathbf{b}_{y}$ denotes a constant that does not depend on $\mathbf{c}$. Together with Eq. (10), we arrive at the following approximation:
$\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y)-\log{Z(\mathbf{x})}\approx p(\mathbf{c}|y)\log p(\mathbf{c}|\mathbf{x})+\mathbf{b}_{y},$ (14)
where the left-hand side represents the logit-based formulation of next-token prediction, and the right-hand side highlights the contribution of the latent variables $\mathbf{c}$. This approximation establishes a connection between the explicit modeling of logits and the probabilistic relationships among the latent variables in the generative framework.
@@
Now, we apply the diversity condition outlined in Section 3, which asserts that there exist $\ell+1$ values of $y$, where $\ell$ is the dimension of $p(\mathbf{c})$ (equivalently $p(\mathbf{c}|y)$ or $p(\mathbf{c})|\mathbf{x}$), i.e., $y_{0}$, $y_{1}$, …, $y_{\ell}$, such that the matrix $\mathbf{L}=\big{(}p(\mathbf{c}|y=y_{1})-p(\mathbf{c}|y=y_{0}),...,p(\mathbf{c}|y=y_{\ell})-p(\mathbf{c}|y=y_{0})\big{)}$ of size $\ell\times\ell$ is invertible. In this context, for $y=0$, we have:
$\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y=0)-\log{Z(\mathbf{x})}\approx p(\mathbf{c}|y=0)\log p(\mathbf{c}|\mathbf{x})+\mathbf{b}_{y=0}.$ (15)
For $y=1$, we similarly obtain::
$\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y=1)-\log{Z(\mathbf{x})}\approx p(\mathbf{c}|y=1)\log p(\mathbf{c}|\mathbf{x})+\mathbf{b}_{y=1},$ (16)
@@
$\big{(}\mathbf{f}_{y}(y=1)-\mathbf{f}_{y}(y=0)\big{)}^{T}\mathbf{f}_{x}(\mathbf{x})\approx\big{(}p(\mathbf{c}|y=1)-p(\mathbf{c}|y=0)\big{)}\log p(\mathbf{c}|\mathbf{x})+(\mathbf{b}_{y=1}-\mathbf{b}_{y=0}).$ (17)
Similar, for $y=2$, we generate the following equation:
$\big{(}\mathbf{f}_{y}(y=2)-\mathbf{f}_{y}(y=0)\big{)}^{T}\mathbf{f}_{x}(\mathbf{x})\approx p(\mathbf{c}|y=2)-p(\mathbf{c}|y=0)\log p(\mathbf{c}|\mathbf{x})+(\mathbf{b}_{y=2}-\mathbf{b}_{y=0}).$ (18)
@@
According to the diversity condition, where $y$ can take $\ell+1$ values, we obtain a total of $\ell$ equations similar to Eqs. (17) and (18). Collecting all of these equations, we have:
$\underbrace{[\mathbf{f}_{y}(y=1)-\mathbf{f}_{y}(y=0)...\mathbf{f}_{y}(y=n+1)-\mathbf{f}_{y}(y=0)]^{T}}_{\mathbf{\hat{L}}^{T}}\mathbf{f}_{x}(\mathbf{x})\approx$ (19)
$\underbrace{[p(\mathbf{c}|y=1)-p(\mathbf{c}|y=0);...p(\mathbf{c}|y=\ell)-p(\mathbf{c}|y=0)]}_{\mathbf{L}}\log p(\mathbf{c}|\mathbf{x})+[\mathbf{b}_{y=1}-\mathbf{b}_{y=0},...,\mathbf{b}_{y=\ell}-\mathbf{b}_{y=0}].$ (20)
@@
According to the diversity condition, the matrix $\mathbf{\hat{L}}=\big{(}\mathbf{f}_{y}(y=y_{1})-\mathbf{f}_{y}(y=y_{0}),...,\mathbf{f}_{y}(y=y_{\ell})-\mathbf{f}_{y}(y=y_{0})\big{)}$ of size $\ell\times\ell$ is invertible, as a result, we arrive:
$\mathbf{f}_{x}(\mathbf{x})\approx\underbrace{(\mathbf{\hat{L}}^{T})^{-1}\mathbf{L}}_{\mathbf{A}}\log p(\mathbf{c}|\mathbf{x})+\mathbf{k}.$ (21)

### v2

@@ 4.3 Unifying the Linear Representation Hypothesis
Corollary 4.2 reveals that the weights $\mathbf{W}$ learned through linear probing are intimately connected to the linear transformation $\mathbf{A}$ from our identifiability results, with the key relationship $\mathbf{W}\mathbf{A}=\mathbf{I}$. This relationship, previously unexplored, provides a unified framework for understanding the Linear Representation Hypothesis across various manifestations. These manifestations include Concepts as Directions, Concept Manipulability, and Linear Probing, all of which are unified through the linear transformation $\mathbf{A}$. Specifically:
  • Concepts as directions and concept manipulability focus on extracting or manipulating the columns of $\mathbf{A}$, which correspond to the latent concepts of interest.
@@
By leveraging the relationship $\mathbf{W}\mathbf{A}=\mathbf{I}$, we demonstrate that these different forms of the linear representation hypothesis are not independent, but are instead intrinsically linked through the same linear transformation $\mathbf{A}$ from our identifiability result. This unified viewpoint underscores how $\mathbf{A}$ serves as the underlying structure that connects concept as directions, concept manipulation, and linear probing.
@@ 5 Experiments on Simulation and Real Data
We begin by conducting experiments on synthetic data, which is generated through the following process: First, we create random directed acyclic graphs (DAGs) with $n$ latent variables, representing concepts. For each random DAG, the conditional probabilities of each variable given its parents are modeled using Bernoulli distributions, where the parameters are sampled uniformly from $[0.2,0.8]$. To simulate a nonlinear mixture process, we then convert the latent variable samples into one-hot format and randomly apply a permutation matrix to the one-hot encoding, generating one-hot observed samples. These are then transformed into binary observed samples. To simulate next-token prediction, we randomly mask part of the binary observed data, i.e., ${x}_{i}$, and use the remaining portion, i.e., $\mathbf{x}_{j}$ where $j\neq i$, to predict the masked part ${x}_{i}$. More details can be found in Section F.1 in Appendix.
@@
Figure 2: (Left) Classification accuracy under varying numbers of observed variables: as the number increases, the latent-to-observed mapping becomes more invertible, leading to improved recovery of latent variables—consistent with our result in Theorem 3.1. (Right) Classification accuracy across different graph structures demonstrates the robustness of the identifiability results to graph structure and latent variable size.
@@
Figure 3: Results of the product $\mathbf{A}_{s}\times\mathbf{W}_{s}$ across the LLaMA-2 and Pythia model families. Here, $\mathbf{A}_{s}$ represents a matrix derived from the feature differences of 27 counterfactual pairs, while $\mathbf{W}_{s}$ is a weight matrix obtained from a linear classifier trained on these features. The product approximates the identity matrix, supporting the theoretical findings outlined in Corollary 4.2. Refer to Appendix H for more results on LLaMA-3 and DeepSeek-R1.
@@
We now present experiments on pre-trained LLMs. While it is challenging to collect all latent variables from real-world data to directly evaluate our identifiability result in Thorem 3.1, we can instead assess our corollaries to indirectly validate it. For Corollary 4.1, prior studies have already demonstrated the linear representation properties of LLM embeddings using counterfactual pairs or linear probing (Mikolov et al., 2013; Pennington et al., 2014; Turner et al., 2023; Li et al., 2024; Wang et al., 2023; Park et al., 2023). Therefore, we shift our focus to a new property in Corollary 4.2, which has not been explored in previous works. To investigate it, we utilize 27 counterfactual pairs from Park et al. (2023), which are constructed based on the Big Analogy Test dataset (Gladkova et al., 2016). More details can be found in Section F.2 in Appendix. Specifically, supported by Corollary 4.1, we compute the differences between the features of these 27 counterfactual pairs on pre-trained LLMs to construct a matrix $\mathbf{A}_{s}$ with dimensions $27\times dim$, where $dim$ corresponds to the feature dimension of the pre-trained LLM used. This matrix can be regarded as a submatrix of a variant of $\mathbf{A}$, in which each row is associated with the probability of a specific latent concept. Additionally, we train a linear classifier using the features of these counterfactual pairs and extract the corresponding weights, supported by Corollary 4.2, forming a weight matrix $\mathbf{W}_{s}$ with dimensions $dim\times 27$, which also serves as a submatrix of $\mathbf{W}$. Finally, we examine the product of $\mathbf{A}_{s}$ and $\mathbf{W}_{s}$ to assess whether it approximates the identity matrix. Figure 3 displays the results of this product across the LLaMA-2 and Pythia model families. The results show that the product approximates the identity matrix, which is consistent with the theoretical finding in Corollary 4.2.
@@ 6 Conclusion and Discussion
Our analysis of LLMs through the lens of latent variable models and identifiability provides valuable insights into the underlying mechanisms that contribute to their success [^1] We strongly encourage the reader to refer to the additional discussions on latent variable models in Appendix G.. We have shown that that the representations learned by LLMs can be approximated as a linear transformation of the posterior distribution of latent variables. This not only reinforces the linear representation hypothesis but also highlights that concepts are encoded in the linear transformation matrix.
@@ Appendix A Related Work
This work is closely related to causal representation learning (Schölkopf et al., 2021), which aims to identify high-level latent causal variables from low-level observational data. Many prior studies (Brehmer et al., 2022; Von Kügelgen et al., 2021; Massidda et al., 2023; von Kügelgen et al., 2023; Ahuja et al., 2023; Seigal et al., 2022; Shen et al., 2022; Liu et al., 2022; Buchholz et al., 2023; Varici et al., 2023; Liu et al., 2024b, 2025, a, c; Hyvarinen & Morioka, 2016; Hyvarinen et al., 2019; Khemakhem et al., 2020) have developed theoretical frameworks supporting the recovery of true latent variables up to simple transformations. However, these works primarily focus on continuous spaces and do not address the next-token prediction framework employed by LLMs. A subset of works (Gu & Dunson, 2023; Kong et al., 2024; Kivva et al., 2021) has explored causal representation learning in discrete spaces for both latent and observed variables, often imposing specific graph structures and assuming invertible mappings from latent to observed spaces. Despite their focus on discrete settings, none of these studies consider the next-token prediction framework in LLMs. In contrast, our work overcomes these limitations. We analyze approximate identifiability without relying on strict invertibility assumptions, which better aligns with the complex and often non-invertible relationships observed in real-world data. While a recent study has examined non-invertible mappings from latent to observed spaces, they rely on additional historical information to effectively restore invertibility. In our approach, we focus on achieving approximate identifiability without requiring such additional information. A very recent work (Rajendran et al., 2024) explores identifiability analysis for LLMs; however, they model both observed text data and latent variables as continuous and still assume invertibility in their analysis. In contrast, our framework explicitly considers discrete latent and observed variables, relaxed invertibility assumptions, and directly aligns with the next-token prediction paradigm, offering a more realistic and generalizable approach to causal representation learning in LLMs.
@@ Appendix B Limitations
Now, let us focus on the right of Eq. (8). Introducing $p(\mathbf{c}|y)$ and $q(\mathbf{c})$, where $q(\mathbf{c})$ is a valid probability distribution such that $\sum_{c}q(\mathbf{c})=1$ , we can obtain that:
$\log p(y|\mathbf{x})$ $=\mathbb{E}_{p(\mathbf{c}|y)}\log p(y|\mathbf{x})$ (10)
$=\mathbb{E}_{p(\mathbf{c}|y)}\log\big{(}p(\mathbf{c}|y,\mathbf{x})\frac{p(y|\mathbf{x})}{p(\mathbf{c}|y,\mathbf{x})}\big{)}$ (11)
$=\mathbb{E}_{p(\mathbf{c}|y)}\log{p(y,\mathbf{c}|\mathbf{x})}-\mathbb{E}_{p(\mathbf{c}|y)}\log{p(\mathbf{c}|y,\mathbf{x})}$ (12)
$=\mathbb{E}_{p(\mathbf{c}|y)}\log{p(\mathbf{c}|\mathbf{x})}+\mathbb{E}_{p(\mathbf{c}|y)}\log{\underbrace{p(y|\mathbf{c},\mathbf{x})}_{p(y|\mathbf{c})}}-\mathbb{E}_{p(\mathbf{c}|y)}\log{p(\mathbf{c}|y,\mathbf{x})}$ (13)
Together with Eq. (7) and Eq. (9), we have:
$\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y)-\log{Z(\mathbf{x})}=\mathbb{E}_{p(\mathbf{c}|y)}\log{p(\mathbf{c}|\mathbf{x})}+\underbrace{\mathbb{E}_{p(\mathbf{c}|y)}\log{{p(y|\mathbf{c})}}}_{b_{y}}-\mathbb{E}_{p(\mathbf{c}|y)}\log{p(\mathbf{c}|y,\mathbf{x})}.$ (14)
Define $\mathbb{E}_{p(\mathbf{c}|y)}\log{{p(y|\mathbf{c})}}=b_{y}$ Define a vector $[p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})]_{\mathbf{c}_{i}}$ as the vector constructed from the probabilities of all possible values of $\mathbf{c}$, conditional on $\mathbf{x}$, where ${\mathbf{c}_{i}\in\mathcal{C}}$ and $\mathcal{C}$ is the set of all possible values that $\mathbf{c}$ can take. As a result, $\mathbb{E}_{p(\mathbf{c}|y)}\log{p(\mathbf{c}|\mathbf{x})}=\sum_{\mathbf{c}}{p(\mathbf{c}|y)}\log{p(\mathbf{c}|\mathbf{x})}=[{p}(\mathbf{c}=\mathbf{c}_{i}|y)]_{\mathbf{c}_{i}}{[\log p(\mathbf{c}|\mathbf{x})]_{\mathbf{c}_{i}}}$. Then we can re-write Eq. (14) as follows:
$\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y)-\log{Z(\mathbf{x})}=$ $[{p(\mathbf{c}=\mathbf{c}_{i}|y)}]^{T}_{\mathbf{c}_{i}}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}$ (15)
$-[{p(\mathbf{c}=\mathbf{c}_{i}|y)}]^{T}_{\mathbf{c}_{i}}[\log{p(\mathbf{c}=\mathbf{c}_{i}|y,\mathbf{x})}]_{\mathbf{c}_{i}}+b_{y}$ (16)
@@
Now, we apply the diversity condition outlined in Section 3, which asserts that there exist $\ell+1$ values of $y$, where $\ell$ is the number of all possible values of $\mathbf{c}$, i.e., $y_{0}$, $y_{1}$, …, $y_{\ell}$, such that the matrix $\mathbf{L}=\big{(}[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{1})]^{T}_{\mathbf{c}_{i}}-[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})]^{T}_{\mathbf{c}_{i}},...,[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{\ell})]^{T}_{\mathbf{c}_{i}}-[p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})]^{T}_{\mathbf{c}_{i}}\big{)}$ of size $\ell\times\ell$ is invertible. In this context, for $y=0$, we have:
$\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y=y_{0})-\log{Z(\mathbf{x})}=$ $[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})}]^{T}_{\mathbf{c}_{i}}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}$
$-\underbrace{[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})}]^{T}_{\mathbf{c}_{i}}[\log{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0},\mathbf{x})}]_{\mathbf{c}_{i}}}_{h_{y_{0}}}+b_{y_{0}},$ (17)
where we define ${h_{y_{0}}}={[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})}]^{T}_{\mathbf{c}_{i}}[\log{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0},\mathbf{x})}]_{\mathbf{c}_{i}}}$. For $y=1$, we similarly obtain:
$\mathbf{f}_{x}(\mathbf{x})^{T}\mathbf{f}_{y}(y=y_{1})-\log{Z(\mathbf{x})}=$ $[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{1})}]^{T}_{\mathbf{c}_{i}}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}$
$-h_{y_{1}}+b_{y_{1}}.$ (18)
@@
$\big{(}\mathbf{f}_{y}(y=y_{1})^{T}-\mathbf{f}_{y}(y=y_{0})^{T}\big{)}\mathbf{f}_{x}(\mathbf{x})=$ $\big{(}[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{1})}]^{T}_{\mathbf{c}_{i}}-[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})}]^{T}_{\mathbf{c}_{i}}\big{)}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}$
$-(h_{y_{1}}-h_{y_{0}})+b_{y_{1}}-b_{y_{0}}.$ (19)
@@
According to the diversity condition, where $y$ can take $\ell+1$ values, we obtain a total of $\ell$ equations similar to Eq. (19). Collecting all of these equations, we have:
$\underbrace{\big{(}\mathbf{f}_{y}(y=y_{1})-\mathbf{f}_{y}(y=y_{0}),...,\mathbf{f}_{y}(y=y_{\ell})-\mathbf{f}_{y}(y=y_{0})\big{)}^{T}}_{\mathbf{\hat{L}}^{T}}\mathbf{f}_{x}(\mathbf{x})$ (20)
$=\underbrace{\big{(}[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{1})}]_{\mathbf{c}_{i}}-[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})}]_{\mathbf{c}_{i}},...,[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{\ell})}]_{\mathbf{c}_{i}}-[{p(\mathbf{c}=\mathbf{c}_{i}|y=y_{0})}]_{\mathbf{c}_{i}}\big{)}^{T}}_{\mathbf{L}}\times$
$[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}-\underbrace{[h_{y_{1}}-h_{y_{0}},...,h_{y_{\ell}}-h_{y_{0}}]}_{\mathbf{h}_{y}}+\underbrace{[b_{y_{1}}-b_{y_{0}},...,b_{y_{\ell}}-b_{y_{0}}]}_{\mathbf{b}_{y}}$ (21)
@@
According to the diversity condition, the matrix $\mathbf{\hat{L}}$ of size $\ell\times\ell$ is invertible, as a result, we arrive:
$\mathbf{f}_{x}(\mathbf{x})=\underbrace{(\mathbf{\hat{L}}^{T})^{-1}\mathbf{L}}_{\mathbf{A}}[\log{p(\mathbf{c}=\mathbf{c}_{i}|\mathbf{x})}]_{\mathbf{c}_{i}}-(\mathbf{\hat{L}}^{T})^{-1}\mathbf{{h}}_{y}+(\mathbf{\hat{L}}^{T})^{-1}\mathbf{b}_{y}.$ (22)